Fill in the missing data for each of the following independent cases. (Ignore income taxes.) (Do Not Found Intermed! calculatPlease show all work
Answer #1

1.

Break even sales = Fixed cost / Contribution margin ratio

83,000 = 66,400/ Contribution margin ratio

Contribution margin ratio = 66,400/83,000

= 80%

Since Contribution margin ratio i 80%, hence variable cost is 20% of sales

Variable cost = Sales x 20%

83,000 = Sales x 20%

Sales = 83,000 / 20%

= $415,000

Contribution margin = Sales - Variable cost

= 415,000-83,000

= $332,000

Net Income = Contribution margin - Fixed cost

= 332,000-66,400

= $265,600

2.

Contribution margin = Sales - Variable cost

33,000 = 150,000-Variable expense

Variable expense = $117,000

Contribution margin ratio = Contribution margin/ Sales revenue

= 33,000/150,000

= 22%

Break even sales = Fixed cost / Contribution margin ratio

150,000 = Fixed cost/ 22%

Fixed cost = $33,000

Net Income = Contribution margin - Fixed cost

= 33,000-33,000

= $0

3.

Contribution margin = Sales - Variable cost

210,000 = Sales - 126,000

Sales = 336,000

Net Income = Contribution margin - Fixed cost

160,000 = 210,000- Fixed cost

Fixed cost = $50,000

Contribution margin ratio = Contribution margin/ Sales revenue

= 210,000/336,000

= 62.5%

Break even sales = Fixed cost / Contribution margin ratio

= 50,000/62.5

= $80,000

4.

Contribution margin = Sales - Variable cost

= 55,000-11,000

= $44,000

Net Income = Contribution margin - Fixed cost

19,300 = 44,000-Fixed cost

Fixed cost = $24,700

Contribution margin ratio = Contribution margin/ Sales revenue

= 44,000/55,000

= 80%

Break even sales = Fixed cost / Contribution margin ratio

= 24,700/80%

= $30,875

Sales revenue Variable expense Total Contribution margin Total Fixed expense Net Income Break-even Sales revenue
1 415,000 83,000 332,000 66,400 265,600 83,000
2 150,000 117,000 33,000 33,000 0 150,000
3 336,000 126,000 210,000 50,000 160,000 80,000
4 55,000 11,000 44,000 24,700 19,300 30,875

Kindly comment if you need further assistance.

Thanks‼!